# DisastIR: A Comprehensive Information Retrieval Benchmark for Disaster Management

### v1

@@ Appendix I Additional Analyses of Model Performance across 48 Tasks
Biological 68.22 59.69 43.28 56.01 36.94 67.79 55.32
Chemical 70.28 59.30 38.93 56.44 33.85 68.50 54.55
Environmental 67.62 57.23 41.34 52.00 38.74 66.61 53.92
Extraterrestrial 65.28 56.57 42.45 51.24 32.11 68.85 52.75
Geohazard 65.65 52.61 35.16 49.45 30.81 64.75 49.74
Meteorological&hydrological 67.20 59.98 38.81 51.14 33.15 66.41 52.78
Societal 69.23 62.26 41.72 53.73 38.01 68.27 55.54
Technological 68.15 60.74 40.77 54.70 34.90 67.23 54.41
Avg. 67.70 58.55 40.31 53.09 34.81 67.30 53.63
@@
Biological 51.58 53.85 50.86 58.41 37.78 65.45 52.99
Chemical 47.97 50.77 48.75 58.64 34.71 67.10 51.32
Environmental 50.63 51.44 47.75 60.93 41.00 63.61 52.56
Extraterrestrial 50.50 57.93 45.21 59.49 33.18 65.59 51.98
Geohazard 47.81 52.76 44.28 56.91 33.09 62.52 49.56
Meteorological&hydrological 49.27 53.06 47.10 59.05 34.13 62.09 50.78
Societal 50.00 59.17 45.32 55.97 36.82 62.81 51.68
Technological 49.56 52.77 46.99 58.39 33.90 66.24 51.31
Avg. 49.67 53.97 47.03 58.47 35.58 64.43 51.52
@@
Biological 56.35 55.10 36.71 55.50 35.62 66.83 51.02
Chemical 56.24 51.49 33.57 56.73 34.65 66.02 49.78
Environmental 57.20 51.53 33.76 54.54 38.85 65.74 50.27
Extraterrestrial 58.77 57.02 33.03 58.30 32.68 66.38 51.03
Geohazard 55.15 53.62 30.65 53.17 35.16 63.51 48.54
Meteorological&hydrological 56.45 54.24 31.36 53.73 33.20 62.50 48.58
Societal 58.61 58.52 32.33 54.25 35.96 63.75 50.57
Technological 57.39 56.25 30.28 55.56 35.44 66.37 50.21
Avg. 57.02 54.72 32.71 55.22 35.19 65.14 50.00
@@
Biological 47.81 47.68 31.38 53.01 28.93 60.77 44.93
Chemical 45.25 42.32 30.68 55.45 27.44 57.61 43.12
Environmental 51.27 45.90 28.36 52.48 31.09 53.90 43.83
Extraterrestrial 48.32 45.50 26.64 53.35 25.57 58.43 42.97
Geohazard 46.74 42.91 25.79 50.88 25.91 55.30 41.25
Meteorological&hydrological 46.89 46.96 26.52 52.61 28.11 55.46 42.76
Societal 47.90 47.18 26.05 45.50 25.78 54.76 41.20
Technological 44.82 45.61 27.12 53.74 24.75 58.68 42.45
Avg. 47.38 45.51 27.82 52.13 27.20 56.86 42.82
@@
Biological 35.16 31.88 27.70 43.16 28.81 50.09 36.13
Chemical 39.05 28.91 27.52 44.70 20.32 39.50 33.33
Environmental 42.55 33.30 30.97 36.88 30.97 50.49 37.53
Extraterrestrial 33.98 33.26 25.64 44.49 28.29 42.68 34.72
Geohazard 35.77 28.05 23.63 33.08 27.89 47.75 32.70
Meteorological&hydrological 33.17 28.16 28.04 32.61 28.43 44.46 32.48
Societal 35.09 36.96 23.60 38.97 30.61 39.84 34.18
Technological 30.12 30.78 25.34 41.78 23.06 50.72 33.63
Avg. 35.61 31.41 26.55 39.46 27.30 45.69 34.34
@@
Biological 72.89 67.94 72.41 69.18 51.66 74.32 68.07
Chemical 72.03 66.49 69.62 72.05 49.51 72.65 67.06
Environmental 72.25 66.24 72.43 73.16 52.99 73.20 68.38
Extraterrestrial 70.01 68.71 67.85 70.71 47.73 75.48 66.75
Geohazard 70.77 65.41 69.06 69.21 49.01 70.81 65.71
Meteorological&hydrological 72.44 67.29 70.11 70.49 50.84 73.12 67.38
Societal 72.78 70.13 71.38 69.38 53.92 74.26 68.64
Technological 70.89 68.42 70.48 70.80 51.20 75.08 67.81
Avg. 71.76 67.58 70.42 70.62 50.86 73.61 67.47
@@
Biological 41.34 30.98 16.73 33.28 35.21 54.92 35.41
Chemical 40.77 30.26 16.39 34.55 33.24 56.39 35.27
Environmental 43.56 34.71 17.68 34.42 40.28 56.81 37.91
Extraterrestrial 41.09 27.96 14.48 31.93 32.31 58.93 34.45
Geohazard 36.34 28.20 12.92 33.83 31.08 53.90 32.71
Meteorological&hydrological 37.75 25.59 14.99 33.70 34.90 57.75 34.11
Societal 42.99 35.02 16.49 28.39 36.37 57.69 36.16
Technological 42.70 30.53 12.85 31.51 33.57 58.20 34.89
Avg. 40.82 30.41 15.32 32.70 34.62 56.82 35.11
@@
Biological 53.33 58.87 38.42 60.63 41.22 62.01 52.41
Chemical 56.40 57.57 40.62 61.11 39.01 63.87 53.10
Environmental 57.52 59.91 45.04 62.53 45.87 63.73 55.77
Extraterrestrial 54.74 56.70 36.38 60.14 37.86 62.82 51.44
Geohazard 54.21 57.26 34.76 59.82 39.80 60.19 51.01
Meteorological&hydrological 51.69 59.76 36.28 61.42 40.67 62.49 52.05
Societal 57.78 63.46 41.87 57.77 44.56 64.80 55.04
Technological 57.02 61.94 36.67 59.97 40.49 67.14 53.87
Avg. 55.34 59.43 38.75 60.42 41.19 63.38 53.09
@@
Biological 31.68 13.54 8.09 38.04 39.39 54.55 30.88
Chemical 33.35 14.43 9.76 35.38 36.26 58.17 31.23
Environmental 35.53 14.87 9.18 35.83 43.61 57.19 32.70
Extraterrestrial 33.40 14.31 8.29 35.18 39.67 58.59 31.57
Geohazard 31.14 14.32 6.15 35.78 36.75 54.06 29.70
Meteorological&hydrological 30.96 14.49 9.29 35.75 39.04 56.34 30.98
Societal 35.17 14.34 10.72 31.87 40.57 57.82 31.75
Technological 35.55 13.70 7.63 34.58 36.14 58.32 30.99
Avg. 33.35 14.25 8.64 35.30 38.93 56.88 31.23
@@
Biological 20.41 30.10 17.56 48.74 42.93 63.69 37.24
Chemical 23.39 29.15 20.20 48.55 41.74 65.45 38.08
Environmental 28.64 33.80 21.74 51.30 49.00 65.53 41.67
Extraterrestrial 25.43 24.07 17.35 46.65 41.79 66.04 36.89
Geohazard 26.06 30.31 17.01 46.63 42.02 63.49 37.59
Meteorological&hydrological 22.61 28.98 17.43 48.37 42.51 64.42 37.39
Societal 29.55 35.48 17.74 47.59 42.61 64.06 39.51
Technological 29.77 32.56 17.47 48.12 41.71 66.91 39.42
Avg. 25.73 30.56 18.31 48.24 43.04 64.95 38.47
@@
Biological 60.58 62.53 48.09 57.58 42.68 64.84 56.05
Chemical 61.22 61.01 47.12 59.78 40.78 64.72 55.77
Environmental 63.96 62.03 49.99 60.78 47.75 65.53 58.34
Extraterrestrial 60.83 61.17 46.86 58.68 39.05 65.74 55.39
Geohazard 60.38 60.30 47.90 57.05 43.33 63.02 55.33
Meteorological&hydrological 60.47 63.37 46.95 58.29 40.69 63.79 55.59
Societal 62.72 65.79 47.63 54.61 43.17 66.73 56.78
Technological 61.34 63.79 47.86 57.56 41.24 67.88 56.61
Avg. 61.44 62.50 47.80 58.04 42.34 65.28 56.23
@@
Biological 34.80 26.88 18.93 44.11 36.25 66.25 37.87
Chemical 33.78 28.03 20.01 45.40 35.16 66.72 38.18
Environmental 39.09 30.43 23.33 42.75 43.67 68.17 41.24
Extraterrestrial 35.15 21.60 16.03 41.94 34.80 66.06 35.93
Geohazard 35.78 25.98 16.63 41.53 36.37 63.31 36.60
Meteorological&hydrological 34.13 28.33 16.74 43.00 38.08 66.33 37.77
Societal 37.91 30.80 17.34 39.73 38.71 67.63 38.69
Technological 37.77 27.63 16.34 41.68 35.15 68.29 37.81
Avg. 36.05 27.46 18.17 42.52 37.27 66.59 38.01
@@
Biological 65.94 61.42 50.92 62.06 46.86 71.23 59.74
Chemical 64.39 57.94 52.68 63.52 46.68 71.24 59.41
Environmental 65.87 60.32 47.60 64.84 52.70 73.30 60.77
Extraterrestrial 63.75 58.73 44.83 62.05 47.34 72.81 58.25
Geohazard 63.08 59.78 44.29 60.60 46.28 70.89 57.49
Meteorological&hydrological 65.83 61.04 45.60 62.71 45.98 71.02 58.70
Societal 66.36 65.53 44.21 62.98 50.31 71.93 60.22
Technological 63.95 63.52 47.03 63.49 48.21 73.06 59.88
Avg. 64.90 61.04 47.14 62.78 48.04 71.94 59.31
@@
Biological 76.15 68.85 71.20 68.79 52.90 78.79 69.45
Chemical 74.65 65.12 68.88 70.95 52.85 76.74 68.20
Environmental 73.45 66.91 69.84 70.05 56.45 77.55 69.04
Extraterrestrial 72.61 68.14 66.38 68.78 48.90 79.31 67.35
Geohazard 74.22 65.10 68.63 67.46 51.55 75.86 67.14
Meteorological&hydrological 74.21 69.31 70.86 68.37 52.17 76.76 68.61
Societal 75.49 70.65 69.64 68.07 57.23 78.49 69.93
Technological 73.03 68.48 69.23 68.84 52.73 78.12 68.41
Avg. 74.23 67.82 69.33 68.91 53.10 77.70 68.52
@@
Biological 72.23 65.55 67.42 65.01 53.59 76.99 66.80
Chemical 71.18 63.09 65.04 67.03 51.93 75.97 65.71
Environmental 70.06 63.56 65.57 68.21 57.26 76.07 66.79
Extraterrestrial 67.37 61.83 63.32 65.93 51.12 77.02 64.43
Geohazard 71.49 62.52 65.70 65.56 52.29 75.73 65.55
Meteorological&hydrological 70.52 65.69 67.86 67.18 52.54 75.66 66.58
Societal 71.90 66.62 66.65 66.51 57.99 76.24 67.65
Technological 70.39 66.48 67.29 66.79 54.18 75.83 66.83
Avg. 70.64 64.42 66.11 66.53 53.86 76.19 66.29
@@
Biological 67.85 62.93 58.90 62.08 47.03 74.82 62.27
Chemical 66.55 59.44 58.58 63.27 45.05 74.71 61.27
Environmental 67.45 63.11 60.69 64.54 49.66 74.99 63.41
Extraterrestrial 63.48 63.63 55.96 63.26 43.63 75.01 60.83
Geohazard 65.17 60.34 56.01 60.72 43.17 74.34 59.96
Meteorological&hydrological 66.14 62.64 59.13 62.21 44.88 74.08 61.52
Societal 64.92 67.05 59.00 60.68 46.63 74.02 62.05
Technological 64.58 65.38 58.65 61.44 44.27 75.15 61.58
Avg. 65.77 63.07 58.37 62.28 45.54 74.64 61.61
@@
Biological 59.31 63.11 56.68 62.40 51.02 75.08 61.27
Chemical 59.77 59.78 54.60 63.05 49.49 75.10 60.30
Environmental 62.61 63.25 57.49 64.79 54.51 75.21 62.98
Extraterrestrial 58.42 63.93 54.57 61.76 48.29 75.67 60.44
Geohazard 59.25 62.26 54.94 60.25 50.44 74.02 60.19
Meteorological&hydrological 60.07 63.55 57.59 62.98 49.77 75.12 61.51
Societal 60.04 66.56 56.92 60.85 53.09 74.75 62.04
Technological 61.74 64.95 56.80 62.20 51.07 74.97 61.95
Avg. 60.15 63.42 56.20 62.29 50.96 74.99 61.34
@@
Biological 18.64 19.90 22.40 32.79 22.57 43.06 26.56
Chemical 25.67 19.23 19.34 36.47 16.04 35.54 25.38
Environmental 24.96 17.65 21.97 28.46 22.05 41.30 26.06
Extraterrestrial 21.42 21.71 18.92 34.04 21.21 37.67 25.83
Geohazard 21.45 17.30 16.98 25.16 19.57 41.11 23.59
Meteorological&hydrological 20.67 15.84 20.54 24.92 20.85 37.62 23.41
Societal 21.53 24.83 16.94 31.62 23.46 35.64 25.67
Technological 18.86 19.62 18.73 34.85 17.34 44.61 25.67
Avg. 21.65 19.51 19.48 31.04 20.39 39.57 25.27
@@
Biological 66.38 63.19 61.91 61.38 48.09 75.04 62.67
Chemical 66.99 60.81 61.30 63.53 47.24 75.17 62.51
Environmental 67.61 64.30 62.38 65.73 50.80 75.57 64.40
Extraterrestrial 63.65 60.90 57.98 60.98 45.13 74.96 60.60
Geohazard 65.73 61.10 59.06 60.21 43.82 74.60 60.75
Meteorological&hydrological 66.89 63.88 61.80 62.53 46.50 74.76 62.73
Societal 64.53 65.51 61.41 60.10 47.63 74.69 62.31
Technological 65.19 64.31 61.32 61.88 47.51 73.84 62.34
Avg. 65.87 63.00 60.89 62.04 47.09 74.83 62.29
@@
Biological 68.63 62.63 64.36 60.27 49.91 74.90 63.45
Chemical 66.87 62.82 63.95 61.04 49.49 74.90 63.18
Environmental 66.74 63.98 64.06 62.20 54.65 74.76 64.40
Extraterrestrial 65.81 63.71 60.26 59.62 49.28 75.41 62.35
Geohazard 66.65 62.97 63.71 58.47 49.68 75.10 62.76
Meteorological&hydrological 68.57 63.87 63.41 60.45 50.40 74.53 63.54
Societal 67.61 67.61 64.44 58.97 54.07 74.86 64.59
Technological 67.47 66.39 65.80 61.08 50.71 76.02 64.58
Avg. 67.29 64.25 63.75 60.26 51.02 75.06 63.61
@@
Biological 70.03 63.24 64.67 66.51 50.24 64.07 63.13
Chemical 67.34 63.58 61.41 68.56 48.65 63.33 62.15
Environmental 68.93 64.34 63.38 69.14 50.64 62.96 63.23
Extraterrestrial 66.19 65.08 61.14 68.44 49.74 66.73 62.89
Geohazard 68.16 62.03 62.75 64.45 46.34 62.08 60.97
Meteorological&hydrological 68.59 66.05 63.42 66.52 50.51 64.22 63.22
Societal 69.62 68.27 63.92 66.58 51.19 65.31 64.15
Technological 68.25 66.61 65.26 67.51 47.72 65.81 63.53
Avg. 68.39 64.90 63.24 67.21 49.38 64.31 62.91
@@
llmrails-ember-v1
Biological 64.90 62.60 49.37 60.23 45.66 74.47 59.54
Chemical 65.00 60.86 47.16 62.76 43.98 75.04 59.14
Environmental 65.68 61.87 47.82 62.25 48.17 74.18 59.99
Extraterrestrial 65.52 65.06 44.55 62.47 42.29 74.28 59.03
Geohazard 63.42 60.96 44.08 59.80 43.70 72.78 57.46
Meteorological&hydrological 64.02 64.12 45.55 60.74 42.42 72.11 58.16
Societal 65.11 66.53 43.32 59.41 46.03 74.16 59.09
Technological 64.87 64.73 45.61 60.78 44.23 74.90 59.19
Avg. 64.82 63.34 45.93 61.06 44.56 73.99 58.95

### v2

@@ Appendix I Performance of Evaluated Models
Biological 68.08 59.51 42.33 55.85 37.09 66.90 54.96
Chemical 70.12 59.08 38.44 56.34 33.66 67.67 54.22
Environmental 67.34 56.92 40.80 51.83 39.08 66.24 53.70
Extraterrestrial 65.06 56.49 41.82 51.08 32.19 68.13 52.46
Geohazard 65.43 52.43 34.88 49.19 30.74 63.90 49.43
Meteorological&hydrological 66.95 59.82 38.31 50.95 32.76 65.32 52.35
Societal 68.89 62.11 41.13 53.47 37.61 67.46 55.11
Technological 67.83 60.59 39.96 54.51 35.22 66.47 54.10
Avg. 67.46 58.37 39.71 52.90 34.79 66.51 53.29
@@
Biological 53.05 53.49 50.94 60.72 42.05 69.93 55.03
Chemical 50.87 49.40 49.24 59.74 40.53 70.65 53.40
Environmental 52.48 51.20 46.84 63.06 46.04 68.80 54.74
Extraterrestrial 52.19 56.57 45.08 59.60 39.80 69.14 53.73
Geohazard 49.40 50.32 43.49 58.21 39.90 66.28 51.27
Meteorological&hydrological 51.49 51.47 46.42 60.84 39.81 66.60 52.77
Societal 52.35 58.35 45.70 59.23 43.55 67.96 54.52
Technological 51.40 52.30 46.54 59.64 39.59 69.15 53.10
Avg. 51.65 52.89 46.78 60.13 41.41 68.56 53.57
@@
Biological 56.20 54.80 36.10 55.29 35.67 66.15 50.70
Chemical 56.24 51.57 33.14 56.65 34.44 65.21 49.54
Environmental 57.11 51.28 33.24 54.35 38.65 65.29 49.99
Extraterrestrial 58.73 56.79 32.68 58.04 33.39 65.66 50.88
Geohazard 55.18 53.50 30.47 52.89 35.70 62.73 48.41
Meteorological&hydrological 56.09 54.19 31.05 53.69 33.13 61.63 48.30
Societal 58.30 58.31 32.08 54.01 35.92 63.03 50.28
Technological 57.18 56.05 29.83 55.32 35.08 65.75 49.87
Avg. 56.88 54.56 32.32 55.03 35.25 64.43 49.75
@@
Biological 58.08 52.41 28.12 55.48 35.55 66.01 49.28
Chemical 56.12 48.37 27.78 57.54 32.94 64.88 47.94
Environmental 60.67 51.59 24.29 56.89 37.56 64.23 49.21
Extraterrestrial 56.59 51.14 23.53 56.96 32.45 65.57 47.71
Geohazard 54.58 48.79 22.84 53.02 30.27 63.40 45.48
Meteorological&hydrological 56.82 51.60 24.47 54.81 32.21 62.84 47.13
Societal 58.51 54.46 25.09 51.98 31.96 63.81 47.63
Technological 53.59 51.55 25.36 55.68 30.71 64.93 46.97
Avg. 56.87 51.24 25.19 55.30 32.95 64.46 47.67
@@
Biological 75.63 71.17 66.30 70.33 52.99 71.98 68.07
Chemical 74.96 69.32 63.35 72.40 50.04 70.41 66.75
Environmental 74.33 68.53 65.57 72.02 55.38 71.09 67.82
Extraterrestrial 73.32 71.07 62.56 72.36 51.42 73.09 67.30
Geohazard 73.79 68.30 62.62 68.15 51.33 68.24 65.40
Meteorological&hydrological 74.41 70.76 63.94 70.04 52.57 71.04 67.13
Societal 75.17 74.31 65.18 69.84 54.54 72.31 68.56
Technological 73.53 70.53 64.22 71.01 52.24 72.65 67.36
Avg. 74.40 70.50 64.22 70.77 52.56 71.35 67.30
@@
Biological 72.68 67.70 71.20 69.09 52.40 73.38 67.74
Chemical 71.82 66.39 68.83 71.86 49.62 71.69 66.70
Environmental 72.08 65.97 71.56 72.98 53.07 72.53 68.03
Extraterrestrial 69.81 68.26 66.93 70.59 48.33 74.66 66.43
Geohazard 70.61 65.05 68.54 68.63 49.72 69.86 65.40
Meteorological&hydrological 72.13 67.12 69.66 70.34 51.06 71.88 67.03
Societal 72.47 69.89 70.67 69.04 53.50 73.32 68.15
Technological 70.40 68.35 69.58 70.57 50.94 74.36 67.37
Avg. 71.50 67.34 69.62 70.39 51.08 72.71 67.11
@@
Biological 41.19 30.92 16.44 33.21 34.85 54.28 35.15
Chemical 40.74 30.23 16.05 34.55 33.12 55.60 35.05
Environmental 43.49 34.59 17.50 34.29 40.35 56.38 37.77
Extraterrestrial 41.20 28.02 14.27 31.82 32.23 58.17 34.28
Geohazard 36.39 28.06 12.86 33.70 31.00 53.15 32.53
Meteorological&hydrological 37.65 25.44 14.77 33.56 34.65 56.80 33.81
Societal 42.77 34.94 16.33 28.26 36.31 56.89 35.92
Technological 42.39 30.46 12.67 31.41 33.02 57.57 34.59
Avg. 40.73 30.33 15.11 32.60 34.44 56.11 34.89
@@
Biological 53.19 58.69 37.59 60.60 41.47 61.42 52.16
Chemical 56.34 57.58 40.17 60.94 39.15 63.01 52.86
Environmental 57.44 59.63 44.30 62.25 46.12 63.24 55.50
Extraterrestrial 54.62 56.50 35.98 60.03 38.19 62.14 51.24
Geohazard 54.13 57.02 34.45 59.46 39.61 59.40 50.68
Meteorological&hydrological 51.56 59.75 35.89 61.26 40.85 61.52 51.80
Societal 57.58 63.32 41.46 57.55 44.13 64.00 54.67
Technological 56.77 61.82 36.28 59.77 40.34 66.39 53.56
Avg. 55.20 59.29 38.26 60.23 41.23 62.64 52.81
@@
Biological 31.64 13.53 7.84 37.99 39.26 54.05 30.72
Chemical 33.30 14.33 9.69 35.25 36.09 57.36 31.00
Environmental 35.37 14.95 9.07 35.77 43.12 56.79 32.51
Extraterrestrial 33.43 14.24 8.24 34.96 39.61 57.93 31.40
Geohazard 31.11 14.33 6.01 35.51 36.50 53.24 29.45
Meteorological&hydrological 30.88 14.45 9.19 35.66 38.65 55.44 30.71
Societal 35.04 14.23 10.67 31.79 40.81 57.21 31.62
Technological 35.36 13.69 7.44 34.34 35.96 57.68 30.74
Avg. 33.26 14.22 8.52 35.16 38.75 56.21 31.02
@@
Biological 20.43 30.04 17.22 48.66 43.01 63.04 37.07
Chemical 23.35 29.02 19.92 48.41 41.47 64.66 37.80
Environmental 28.52 33.59 21.38 51.31 48.85 65.02 41.45
Extraterrestrial 25.42 24.03 17.19 46.50 41.84 65.34 36.72
Geohazard 25.97 30.09 16.87 46.30 41.72 62.68 37.27
Meteorological&hydrological 22.59 28.84 17.24 48.22 42.64 63.29 37.14
Societal 29.41 35.38 17.63 47.40 42.67 63.35 39.31
Technological 29.59 32.49 17.24 47.99 41.66 66.24 39.20
Avg. 25.66 30.43 18.09 48.10 42.98 64.20 38.24
@@
Biological 60.41 62.36 47.15 57.49 42.82 64.21 55.74
Chemical 61.13 60.93 46.65 59.61 40.91 64.03 55.54
Environmental 63.92 61.70 49.28 60.57 47.70 65.08 58.04
Extraterrestrial 60.68 60.86 46.24 58.47 39.51 65.04 55.13
Geohazard 60.29 60.11 47.50 56.66 43.53 62.19 55.05
Meteorological&hydrological 60.22 63.23 46.42 58.08 40.62 62.84 55.23
Societal 62.52 65.63 47.14 54.42 43.23 65.96 56.48
Technological 61.07 63.65 47.19 57.40 41.10 67.18 56.27
Avg. 61.28 62.31 47.20 57.84 42.43 64.56 55.94
@@
Biological 38.45 28.71 22.49 49.59 39.71 66.50 40.91
Chemical 36.35 29.04 23.56 49.73 37.62 66.73 40.51
Environmental 41.72 30.95 26.34 48.17 46.21 68.23 43.60
Extraterrestrial 37.42 22.95 18.23 46.11 37.85 66.97 38.26
Geohazard 37.63 26.73 19.34 45.56 38.60 64.43 38.71
Meteorological&hydrological 37.03 29.92 21.05 47.02 39.78 66.11 40.15
Societal 41.05 32.44 21.38 45.08 41.98 67.94 41.65
Technological 39.89 29.84 19.03 47.15 38.40 68.66 40.50
Avg. 38.69 28.82 21.43 47.30 40.02 66.95 40.54
@@
Biological 65.76 61.24 49.90 61.91 46.66 70.40 59.31
Chemical 64.14 57.95 52.14 63.28 46.64 70.31 59.08
Environmental 65.74 60.20 47.01 64.65 52.02 72.71 60.39
Extraterrestrial 63.53 58.39 44.19 61.87 48.12 72.01 58.02
Geohazard 62.93 59.57 43.92 60.19 46.92 69.93 57.25
Meteorological&hydrological 65.40 60.87 45.27 62.50 46.00 69.76 58.30
Societal 66.06 65.24 43.73 62.84 50.83 71.08 59.96
Technological 63.50 63.37 46.26 63.20 47.71 72.31 59.39
Avg. 64.63 60.85 46.55 62.56 48.11 71.06 58.96
@@
Biological 74.92 68.15 68.16 65.94 52.10 77.21 67.75
Chemical 73.23 64.74 66.32 67.34 51.42 75.34 66.40
Environmental 72.43 65.78 65.62 66.67 54.90 76.13 66.92
Extraterrestrial 70.64 66.50 63.10 64.90 47.66 77.47 65.04
Geohazard 73.01 64.51 66.10 64.32 51.30 73.88 65.52
Meteorological&hydrological 72.60 68.24 68.24 65.71 50.53 74.64 66.66
Societal 74.06 69.47 67.21 65.80 55.77 76.62 68.15
Technological 71.80 67.95 66.18 65.39 52.48 76.69 66.75
Avg. 72.84 66.92 66.37 65.76 52.02 76.00 66.65
@@
Biological 71.44 65.78 63.63 64.76 54.11 74.52 65.71
Chemical 69.49 63.79 61.76 65.78 51.92 73.97 64.45
Environmental 68.35 62.30 63.38 67.12 56.62 74.42 65.36
Extraterrestrial 66.65 61.14 60.58 64.80 51.64 75.18 63.33
Geohazard 70.18 62.68 62.51 63.96 53.24 73.25 64.31
Meteorological&hydrological 69.75 65.58 64.54 66.51 53.45 72.88 65.45
Societal 70.68 66.83 64.26 65.48 58.05 74.21 66.59
Technological 69.22 67.09 64.00 65.52 54.13 73.28 65.54
Avg. 69.47 64.40 63.08 65.49 54.14 73.96 65.09
@@
Biological 67.77 62.64 57.97 62.07 46.84 73.89 61.86
Chemical 66.40 59.51 57.92 63.12 45.14 73.75 60.98
Environmental 67.34 62.93 59.94 64.35 49.39 74.33 63.05
Extraterrestrial 63.25 63.36 55.28 63.11 44.20 74.21 60.57
Geohazard 65.05 60.12 55.61 60.40 43.29 73.40 59.64
Meteorological&hydrological 65.83 62.57 58.66 62.12 44.30 72.80 61.04
Societal 64.62 66.76 58.51 60.48 46.61 73.13 61.68
Technological 64.06 65.39 58.19 61.24 44.36 74.32 61.26
Avg. 65.54 62.91 57.76 62.11 45.52 73.73 61.26
@@
Biological 59.28 62.94 55.52 62.28 50.83 74.14 60.83
Chemical 59.63 59.82 53.97 62.90 49.44 74.20 59.99
Environmental 62.62 62.69 56.69 64.54 54.63 74.58 62.62
Extraterrestrial 58.19 63.70 53.88 61.32 49.10 74.81 60.17
Geohazard 59.19 62.11 54.49 59.83 50.74 73.16 59.92
Meteorological&hydrological 59.85 63.30 56.85 62.79 49.28 73.84 60.98
Societal 59.99 66.40 56.31 60.57 53.01 73.80 61.68
Technological 61.49 64.96 56.12 62.00 50.60 74.21 61.56
Avg. 60.03 63.24 55.48 62.03 50.96 74.09 60.97
@@
Biological 66.33 64.74 65.22 67.64 49.45 66.61 63.33
Chemical 66.09 63.90 62.22 69.82 45.50 65.29 62.13
Environmental 66.78 63.40 64.84 69.87 49.74 66.29 63.49
Extraterrestrial 63.51 65.90 61.75 67.95 45.71 68.69 62.25
Geohazard 65.17 63.03 61.52 66.19 45.90 63.45 60.88
Meteorological&hydrological 66.28 65.67 63.71 67.20 47.76 65.58 62.70
Societal 66.48 68.75 64.40 66.21 49.86 67.39 63.85
Technological 64.58 65.91 63.73 68.63 47.54 67.85 63.04
Avg. 65.65 65.16 63.42 67.94 47.68 66.39 62.71
@@
Biological 66.22 63.00 60.72 61.14 48.06 74.11 62.21
Chemical 66.80 60.74 60.79 63.40 47.30 74.23 62.21
Environmental 67.56 63.90 61.53 65.49 50.76 74.97 64.04
Extraterrestrial 63.49 60.68 57.15 60.75 45.70 74.18 60.33
Geohazard 65.60 60.94 58.51 59.69 44.13 73.69 60.42
Meteorological&hydrological 66.63 63.81 61.09 62.35 46.44 73.46 62.30
Societal 64.22 65.40 60.63 59.83 47.62 73.70 61.90
Technological 64.74 64.24 60.35 61.59 46.94 73.10 61.83
Avg. 65.66 62.84 60.10 61.78 47.12 73.93 61.90
@@
Biological 68.44 62.41 63.18 60.18 49.79 73.97 63.00
Chemical 66.63 62.89 63.42 60.94 49.54 74.04 62.91
Environmental 66.66 63.64 63.14 62.10 54.46 74.08 64.01
Extraterrestrial 65.53 63.50 59.49 59.41 50.22 74.51 62.11
Geohazard 66.68 62.72 63.21 58.09 50.76 74.09 62.59
Meteorological&hydrological 68.26 63.78 62.80 60.25 50.22 73.31 63.10
Societal 67.34 67.44 63.71 58.75 54.31 74.00 64.26
Technological 67.13 66.28 64.96 60.76 50.31 75.16 64.10
Avg. 67.08 64.08 62.99 60.06 51.20 74.14 63.26
@@
Biological 69.67 62.94 63.35 66.46 50.04 63.39 62.64
Chemical 67.29 63.69 60.79 68.42 48.02 62.54 61.79
Environmental 68.82 63.96 62.62 68.88 49.87 62.37 62.76
Extraterrestrial 66.07 64.72 60.45 68.28 49.35 66.02 62.48
Geohazard 67.91 61.88 62.11 63.99 45.83 61.20 60.49
Meteorological&hydrological 68.24 65.91 62.79 66.38 49.79 63.12 62.70
Societal 69.27 68.18 63.13 66.23 50.28 64.50 63.60
Technological 67.85 66.46 64.43 67.08 46.77 65.11 62.95
Avg. 68.14 64.72 62.46 66.96 48.75 63.53 62.43
@@
snowflake-arctic-embed-m-long
Biological 19.40 10.57 21.85 39.39 41.02 52.71 30.82
Chemical 19.97 10.58 21.06 36.59 40.40 55.39 30.66
Environmental 21.57 11.81 21.01 34.99 48.55 55.06 32.16
Extraterrestrial 21.22 11.89 17.32 35.89 41.66 56.66 30.77
Geohazard 21.78 8.92 17.99 35.32 40.09 52.50 29.44
Meteorological&hydrological 19.71 10.04 18.89 36.12 40.67 54.56 30.00
Societal 23.30 11.95 20.69 35.64 43.05 56.58 31.87
Technological 24.50 10.93 17.08 35.66 39.77 56.53 30.75
Avg. 21.43 10.84 19.49 36.20 41.90 55.00 30.81